the inlet takes 39 hrs to fill. the drain takes 40 hrs to empty. If the pool is half full and then the inlet and drain are opened. how long from that time will it take for the pool to fill?
Answer by lwsshak3(11628)

the inlet takes 39 hrs to fill. the drain takes 40 hrs to empty. If the pool is half full and then the inlet and drain are opened. how long from that time will it take for the pool to fill?
**
1/39=inlet flow-rate
1/40=drain flow-rate
x=hrs to fill the half-full pool
..
(1/39)*x-(1/40)*x=50%=1/2
x/39-x/40=1/2
(40x-39x)/40*39=1/2
x/1560=1/2
2x=1560
x=780 hrs
ans:
It will take 780 hrs to fill the half-full pool with the inlet and drain opened.